OpenStudy (anonymous):

Please check my answers: 1. Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal Givens: Mass of metal (AI) = 27.776g Distilled water measurement = 26mL Distilled water temperature = 25.4 C Temperature of metal = 100.8 C Temperature of mixture (after metal is placed in calorimeter with distilled water) = 28.9 C

OpenStudy (anonymous):

2. Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the metal in your calculation.

OpenStudy (anonymous):

My answer for question 1 is: Qwater = 27.776 * 4.18 J / (g * °C) * (25.4 * 38.9) = 6562.388 calories or 27475.410 Joules and for question 2 it's: Qmetal = 27.776 * c * (100.8 * 38.9) = -3.816 cal / g * C or -15.980 j / g * C

OpenStudy (anonymous):

@gerryliyana @leonardo0430 @vera_ewing @uri @CausticSyndicalist @robtobey

OpenStudy (joannablackwelder):

Ok, good try, but there are a few errors. One error is that qwater=-qmetal, since the water absorbs the heat that the metal gives off.

OpenStudy (joannablackwelder):

Another error is that when we are calculating qwater, we need to use the mass of the water, not the mass of the metal.

OpenStudy (joannablackwelder):

And another error is that delta T is T2-T1, not times.

OpenStudy (anonymous):

Thank you!! @JoannaBlackwelder